A contractor is considering a sale that promises a profit of $37000 with a probability of 0.7 or a loss(due to bad weather, strikes, and such) if $18000 with a probability of 0.3. What is the expected profit?
Answer by fractalier(2101)

(37000)(.7) + (-18000)(.3) = 25900 - 5400 = $20500